# FORGE: Force-Guided Exploration for Robust Contact-Rich Manipulation under Uncertainty

### v1

@@ V-E Multi-Stage Assembly
We ran $5$ trials resulting in the following success rates: Ring Insertion ($/5$ 5), Small Gear ($/15$ 15), Large Gear ($3/5$), M16 Nut ($/15$ 15). Early terminations saved on average $65s$ in a single trial compared to executing policies for a fixed duration. Overall, the complete assembly succeeded in $3/5$ trials where the failures correspond to the large gear insertion (which has to align the teeth of three already inserted small gears). Please see the accompanying video for a demonstration of the multi-stage assembly and Appendix -C for more experimental details.
@@ VI Related Work
Recent advances in contact-rich simulation have enabled efficient simulation of assembly tasks [5, 6, 52, 53]. However, as discussed throughout the paper, the key challenge becomes the sim-to-real gap: how can we safely and successfully deploy policies that were trained in simulation?
@@
Most prior works focus on insertion-style tasks. We show how the combined application of a force-threshold and dynamics randomization can lead to robust sim-to-real transfer for a range of tasks, including the complicated nut-threading task. Prior work on sim-to-real for nut-threading [14] focuses on large parts ($M48$ nuts) that were fixed to the gripper. In addition, we show these techniques are applicable for sim-to-real transfer of early termination procedures.
Early-Termination: Previous sim-to-real approaches execute policies for a fixed duration [7]. Instead, we would like to terminate once success is achieved. For some tasks, success can be manually specified from sensor data [60, 61]. For others, a classifier can be learned from visual data [62, 63]. However, for contact-rich tasks, visual and proprioceptive data alone may be insufficient to determine success [64]. In such cases, the robot can execute actions to verify success [65]. Previous work learns a separate policy to check success after task execution [24]. Instead, we jointly trained a policy to predict success during task execution.
@@ VII Conclusion
In conclusion, we present FORGE, a method to train sim-to-real policies for robust execution with pose estimation uncertainty. FORGE uses a force threshold and dynamics randomization to learn safe exploration behaviours, enabling successful policy execution with up to $5mm$ of position estimation error. In addition, FORGE can predict early termination, allowing efficient policy execution. In future work, we plan to investigate torque sensing to help develop more efficient search strategies. We also believe research in real-to-sim will help automatically tune simulation models for robust transfer in complicated tasks such as nut-threading.
@@ ACKNOWLEDGMENT
The authors thank Chad Kessens, the Seattle Robotics Lab, and the Robust Robotics Group for their valuable feedback.

### v2

@@ V-E Multi-Stage Assembly
We ran $5$ trials resulting in the following success rates: Ring Insertion ($/5$ 5), Small Gear ($/15$ 15), Large Gear ($3/5$), M16 Nut ($/15$ 15). Early terminations saved on average $65s$ in a single trial compared to executing policies for a fixed duration. Overall, the complete assembly succeeded in $3/5$ trials where the failures correspond to the large gear insertion (which has to align the teeth of three already inserted small gears). Please see the accompanying video for a demonstration of the multi-stage assembly and App. -F for more experimental details.
@@ VI Related Work
Recent advances in contact-rich simulation has enabled efficient simulation of assembly tasks [54, 55, 5, 6]. However, as discussed throughout the paper, the key challenge becomes the sim-to-real gap: how can we safely and successfully deploy policies that were trained in simulation?
@@
Most prior work focus on insertion-style tasks. We show how the combined application of a force-threshold and dynamics randomization can lead to robust sim-to-real transfer for a range of tasks, including the complicated nut-threading task. Prior work on sim-to-real for nut-threading [17] focused on large parts ($M48$ nuts) that were fixed to the gripper. In addition, we show these techniques are applicable for sim-to-real transfer of success prediction.
Success Prediction: Previous sim-to-real approaches execute policies for a fixed duration [7]. Instead, we would like to terminate once success is achieved. For some tasks, success can be manually specified from sensor data [59, 60]. For others, a classifier can be learned from visual data [61, 62]. However, for contact-rich tasks, visual and proprioceptive data alone may be insufficient to determine success [63]. In such cases, the robot can execute actions to verify success [64]. Previous work learns a separate policy to check success after task execution [16]. Instead, we jointly trained a policy to predict success during task execution.
@@ VII Conclusion
In conclusion, we present FORGE, a force-aware method to train robust sim-to-real policies with pose estimation uncertainty. FORGE uses a force threshold and dynamics randomization to learn safe exploration behaviours, enabling successful policy execution with up to $5mm$ of position estimation error. In addition, FORGE can predict task success, allowing efficient policy execution and force threshold tuning. In future work, we plan to investigate torque sensing for more efficient search strategies. We also believe research in real-to-sim will help automatically tune simulation models for more adaptive behaviours.
@@ ACKNOWLEDGMENT
The authors thank the Seattle Robotics Lab and the Robust Robotics Group for their valuable feedback.